# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v2

@@ A.13 Discussion: Relation between load balancing and previous results
A.14 Methods: Analysis of translational limits & gene-specific load analysis
@@
To analyze the affect of increased protein load, we modify Eq. 18:
$\ln\frac{k}{k_{0}}=\textstyle-(\Lambda+\lambda_{p}E)\mu_{m}-\frac{1}{\ln 2}\gamma(\frac{\mu_{m}}{\ln 2},\frac{\phi_{p}}{E\ln 2}),$ (58)
to contain the supplemental load factor $\lambda_{p}$ which is unity if the only protein load is metabolic and $\lambda_{p}>1$ if there is additional load (e.g. toxicity). The optimization conditions (Eqs. 19 and 20) become:
@@
$\textstyle\Lambda\ln 2$ $=$ $-\partial_{\hat{\mu}_{m}}\gamma(\frac{\hat{\mu}_{m}}{\ln 2},\frac{\hat{\mu}_{m}}{\hat{o}\ln 2}),$ (61)
@@
A.15 Results: Increased protein-specific cost reduces the optimal translation efficiency.
@@
Returning to our original motivation, we can understand how genes with a higher protein-to-message cost migrate downwards and rightwards off the optimal $\lambda_{p}=1$ curve, predicting a cloud versus a narrow strip in proteome fraction measurements shown in Fig. S7. If the relative load $\Lambda$ were directly measured, we would expect the predicted optimal translation efficiency curve for $\lambda_{p}=1$ to lie at the top edge of the observed data cloud rather than the bisecting it. This bisection is the consequence of fitting an effective relative load parameter to the abundance data in the unaugmented RLTO model.
A.16 Discussion: Translation limits in E. coli
A critical assumption in the RLTO model to this point has been that the optimal central dogma parameters are realizable in the cell; however, translation can be limited by a number of different mechanisms. The superior performance of the constant- over the optimal-translation-efficiency model in E. coli (Fig. S7D) suggests that this assumption may not be satisfied for bacteria. How do translation limits affect the model phenomenology?
@@
A.17 Methods: Estimate of the message cost and metabolic load
@@
The optimal translation efficiency is (Eq. 23):
@@
and therefore the optimal protein number is:
@@
We define the normalization constant $A$:
$A=\textstyle\sum_{i}[\hat{\mu}_{m}\cdot\frac{1}{\Lambda\ln 2}p_{\Gamma}(\hat{\mu}_{m}|\frac{\hat{\mu}_{m}}{\ln 2},\hat{o}\ln 2)]_{i},$ (69)
where the index $i$ runs over all genes. Now, by summing Eq. 68, over all genes, we derive an expression for the total protein number $N_{p}^{\rm tot}$ in terms of the message cost $\lambda$ and the normalization constant $A$:
@@
This message cost estimate $\hat{\lambda}$ can then be plugged into the metabolic load definition (Eq. 17) to estimate the multiplicity:
@@
A.17.1 Detailed protocol
We first estimate the message numbers, as described in Sec. C.2.3, from data. For each gene $i$, we set the optimal message number equal to the observed message number and then compute the optimal overabundance from the message number using Eq. 26. (Since the result is independent of the assumed $\Lambda$ value, we set an arbitrary initial value of $\Lambda=10^{-5}$.) We then use these single gene optimal message number and overabundances to compute $A$ using Eq. 69. In Eqs. 71 and 72, we use the $N^{\rm tot}_{p}$ from Tab. S2. $N_{m}^{\rm tot}$ is computed by summing the estimated message numbers.
A.17.2 Estimate the message cost and metabolic load in yeast
@@
A.17.3 Estimate the message cost and metabolic load in human cells
@@
A.17.4 Estimate of the modified relative load in bacterial cells
@@
$N_{0}$ $=$ $\sum_{i}(\lambda+\varepsilon){\mu_{m}}=(\lambda+\varepsilon)N_{m},$ (81)
@@ B.2 Results: Overabundance is a robust prediction
Model organism
Growth condition
Doubling time:
Message lifetime:
Message recycling ratio:
messages /cell:
messages /cell-cycle:
proteins:
translation efficiency:
translation rate:
$T$
$\tau_{m}=\gamma_{m}^{-1}$
$m=T/\tau_{m}$
$N^{\rm tot}_{m/c}$
$N^{\rm tot}_{m}$
$N^{\rm tot}_{p}$
$\varepsilon$
[…]
Escherichia coli
LB
30 min [12]
$2.5$ min [13]
12
$7.8\times 10^{3}$ [14]
$9.4\times 10^{4}$
$3\times 10^{6}$ [15]
$22$
$530$
(E. coli)
M9
90 min [12]
$2.5$ min [13]
36
$2.4\times 10^{3}$ [14]
$8.6\times 10^{4}$
$3\times 10^{6}$ [15]
$24$
$580$
Sacchromyces cerevisiae (Yeast–haploid)
YEPD
90 min [16]
22 min [17]
4
$2.9\times 10^{4}$ [18]
$1\times 10^{5}$
$5\times 10^{7}$ [19]
$4\times 10^{2}$
$410$
Mus musculus (Mammalian-mouse)
Tissue
27.5 h [9]
15 h [9]
1.8
$1.7\times 10^{5}$ [9]
$3\times 10^{5}$ [9]
$3\times 10^{9}$ [9]
$1\times 10^{4}$
$660$
Homo sapiens (Human)
Tissue
24 h [20]
14 h [21]
1.7
$3.6\times 10^{5}$ [16]
$5\times 10^{5}$
$2\times 10^{9}$ [15]
$4\times 10^{3}$
120
@@ C.1 Methods: Selection of central dogma parameter estimates
Essential gene classification. The classification of essential genes in yeast comes from the construction of the Keio knockout collection from Baba et al. [23].
@@ C.1.3 Human data
Figure S10: Panel A: One-message-rule for essential genes. For highly transcription genes (high $\mu_{m}$), little compensation for noise is required and the optimal message number tracks with the threshold message number $n_{m}$. However, as the threshold message number approaches one ($n_{m}\rightarrow 1$), the noise is comparable to the mean, and the optimal message number $\mu_{m}$ increases to compensate for the noise. As a result, a lower threshold of roughly one message per cell-cycle is required for essential genes. This threshold is predicted for both fixed (dashed) and optimized translation efficiency (solid). The threshold is weakly dependent on relative load $\Lambda$. Panel B: A one message threshold is observed in three evolutionarily-divergent organisms. As predicted by the RLTO model, essential, but not nonessential genes, are observed to be expressed above a one message per cell-cycle threshold. All organisms have roughly similar distributions of message number for essential genes, which are not observed for message numbers below a couple per cell cycle. Panel C: The distribution of gene transcription rate. The typical transcription rate varies by two orders-of-magnitude between organisms. Panel D: The distribution of gene cellular message number. There is also a two-order-of-magnitude variation between typical cellular message numbers. No consistent lower threshold is observed for either statistic.
@@ C.2.2 Estimating the transcription rate: $\beta_{m}$
To estimate the transcription rate for gene $i$, we start from the estimated cellular message number $N_{m/c,i}$ and use the telegraph model prediction for the cellular message number:
@@ C.2.3 Estimating the message number: $\mu_{m}$
To estimate the message number of gene $i$, we use the predicted value from the telegraph model:
@@ C.3 Results: Histograms of central dogma transcriptional statistics
Gene Message
Annotated function
Essential (E)/
name number:
from Ecocyc
Nonessential (N)
@@
alsK 0.3
The alsK gene encodes a D-allose kinase. Its role in the degradation of D-allose is unclear; AlsK is not required for utilization of a D-allose carbon source; this effect may be due to the presence of other ambiguous sugar kinases within E. coli K-12.
E, N, N
bcsB 0.4
BcsB is encoded in a predicted operon together with bcsA, bcsZ and bcsC. In other organisms, these genes are involved in cellulose biosynthesis, a characteristic of the rdar (red, dry and rough) morphotype. However, the K-12 laboratory strain of E. coli does not show a rdar morphotype and does not produce cellulose.
E, N, N
entD 0.4
AcpS is the founding member of a 4’-phosphopantetheinyl (P-pant) transferase protein family that includes E. coli EntD, E. coli o195 protein, and Bacillus subtilis Sfp; family members share two conserved motifs but relatively low sequence identity overall.
E, N, N
yafF 0.4
No information about this protein was found by a literature search conducted on April 19, 2017.
E,-, N
yagG 0.6
yagGH is predicted to be a member of the XylR regulon; its products may mediate transport (YagG) and hydrolysis (YagH) of xylooligosaccharides; putative XylR and CRP binding sites are identified upstream of yagGH.
E,-, N
yceQ 0.2
No information about this protein was found by a literature search conducted on July 12, 2017.
E, E, N
ydiL 0.2
No information about this protein was found by a literature search conducted on April 7, 2017.
E, N, N
yhhQ 0.4
YhhQ is an inner membrane protein implicated in the uptake of queuosine (Q) precursors - 7-cyano-7-deazaguanine (preQ0) and 7-aminomethyl-7-deazaguanine (preQ1) - for Q salvage. Q-modified tRNA is absent in $\Delta$queD and $\Delta$queD $\Delta$yhhQ strains grown in minimal media with glycerol; Q-modified tRNA is detected when a $\Delta$queD strain is grown in minimal media plus 10 nM preQ0 or preQ1 but is absent when a $\Delta$queD $\Delta$yhhQ strain is grown under these conditions. yhhQ expressed from a plasmid restores the presence of Q-modified tRNA in a $\Delta$queD $\Delta$yhhQ strain.
E,-, N
yibJ 0.3
No information about this protein was found by a literature search conducted on July 9, 2018.
E, N, N
ymfK 0.4
YmfK is a component of the relic lambdoid prophage e14 and is likely the SOS-sensitive repressor. It is similar to the P34 gene of the Shigella flexneri bacteriophage SfV and belongs to the LexA group of SOS-response transcriptional repressors.
E, E, E
@@ Appendix D Analysis of gene-expression noise
This section provides a detailed development of gene expression noise. Sec. D.1 provides a self-contained development of the noise models developed by others which are the input to the RLTO model. Secs. D.2-D.7 describe the RLTO prediction of non-canonical noise scaling and the test of this model.
D.1 Methods: Detailed description of the noise model
D.1.1 Telegraph model: Stochastic kinetic model for central dogma.
The telegraph model for the central dogma describes multiple steps in the gene expression process: Transcription generates mRNA messages [38]. These messages are then translated to synthesize the protein gene products [38]. Both mRNA and protein are subject to degradation and dilution [39]. At the single cell level, each of these processes are stochastic. We will model these processes with the stochastic kinetic scheme [38]:
[…]
D.1.2 Statistical model for protein abundance.
To study the stochastic dynamics of gene expression, we used a stochastic Gillespie simulation [42, 43]. (See Sec. D.1.3.) In particular, we were interested in the explicit relation between the kinetic parameters $(\beta_{m},\gamma_{m},\beta_{p},\gamma_{p})$ and experimental observables.
Consistent with previous reports [44, 45], we find that the distribution of protein number per cell (at cell birth) was described by a gamma distribution: $N_{p}\sim\Gamma(a,\theta)$, where $N_{p}$ is the protein number at cell birth and $\Gamma$ is the gamma distribution which is parameterized by a scale parameter $\theta$ and a shape parameter $a$. (See Sec. D.1.4.) The relation between the four kinetic parameters and these two statistical parameters has already been reported, and have clear biological interpretations [45]: The scale parameter:
$\theta=\varepsilon\ln 2,$ (94)
is proportional to the translation efficiency:
$\varepsilon\equiv\textstyle\frac{\beta_{p}}{\gamma_{m}},$ (95)
$a=\textstyle\frac{\beta_{m}}{\gamma_{p}};$ (96)
$\mu_{m}\equiv\beta_{m}T=a\ln 2,$ (97)
D.1.3 Gillespie simulation of the telegraph model
[…]
The gamma distributions for each mean message number with scale and shape parameters determined by the corresponding translation efficiency and message number ($\theta=\varepsilon\ln 2$, $a=\frac{\mu_{m}}{\ln 2}$) as used for the Gillespie simulation were also plotted with the protein distributions. We observe an excellent match between these Gillespie simulations and the statistical noise model (i.e. gamma function) as shown in Fig. S11.
D.1.4 Gamma function and distribution conventions
$X\sim\Gamma(a,\theta),$ (98)
where $\Gamma(a)$ is the gamma function. The CDF is therefore:
$=$ $P_{\Gamma}(\textstyle\frac{x}{\theta}|a,1),$ (101)
$=$ $\gamma(a,x/\theta),$ (103)
where $\gamma$ is the regularized incomplete gamma function.
D.2 Results: RLTO model predicts non-canonical noise scaling
@@
In E. coli, the translation efficiency is roughly constant (i.e.  $\hat{\mu}_{p}\propto\hat{\mu}_{m}$ , Fig. S7D) and therefore Eq. 106 is consistent with the canonical noise model (Eq. 104). However, in eukaryotes, the translation efficiency grows with message number (i.e.  $\hat{\mu}_{p}\propto\hat{\mu}_{m}^{2}$ , Fig. S7BC). If we substitute this proportionality into Eq. 106, we predict the non-canonical noise scaling:
@@
D.3 Methods: Analysis of gene expression noise

### v3

@@ A.14 Discussion: Relation between load balancing and previous results
A.15 Methods: Analysis of translational limits & gene-specific load analysis
@@
To analyze the effect of increased protein load, we modify Eq. 54:
$\ln\frac{k}{k_{0}}=\textstyle-(\Lambda+\lambda_{p}E)\mu_{m}-\frac{1}{\ln 2}\gamma(\frac{\mu_{m}}{\ln 2},\frac{n_{p}}{\varepsilon\ln 2}),$ (94)
to contain the supplemental load factor $\lambda_{p}$ which is unity if the only protein load is metabolic and $\lambda_{p}>1$ if there is additional load (e.g. toxicity). The optimization conditions (Eqs. 55 and 56) become:
@@
$\textstyle\Lambda\ln 2$ $=$ $\textstyle-\partial_{\hat{\mu}_{m}}\gamma(\frac{\hat{\mu}_{m}}{\ln 2},\frac{\hat{\mu}_{m}}{\hat{o}\ln 2}),$ (97)
@@
A.16 Results: Increased protein-specific cost reduces the optimal translation efficiency.
@@
Returning to our original motivation, we can understand how genes with a higher protein-to-message cost migrate downwards and rightwards off the optimal $\lambda_{p}=1$ curve, predicting a cloud versus a narrow strip in proteome fraction measurements shown in Fig. 4. If the relative load $\Lambda$ were directly measured, we would expect the predicted optimal translation efficiency curve for $\lambda_{p}=1$ to lie at the top edge of the observed data cloud rather than the bisecting it. This bisection is the consequence of fitting an effective relative load parameter to the abundance data in the unaugmented RLTO model.
A.17 Discussion: Translation limits in E. coli
A critical assumption in the RLTO model to this point has been that the optimal central dogma parameters are realizable in the cell; however, translation can be limited by a number of different mechanisms. The superior performance of the constant- over the optimal-translation-efficiency model in E. coli (Fig. 4D) suggests that this assumption may not be satisfied for bacteria. How do translation limits affect the model phenomenology?
@@
A.18 Methods: Estimate of the message cost and metabolic load
@@
The optimal translation efficiency for gene $i$ is (Eq. 59):
@@
and therefore the optimal protein number for gene $i$ is:
@@
We define the normalization constant $A$ (where we restore the explicit gene $i$ subscript):
$A=\textstyle\sum_{i}\hat{\mu}_{m,i}\cdot\frac{1}{\Lambda\ln 2}p_{\Gamma}(\hat{\mu}_{m,i}|\frac{\hat{\mu}_{m,i}}{\ln 2},\hat{o}_{i}\ln 2),$ (105)
where we have restored the explicit gene $i$ index running over all genes. Now, by summing Eq. 104, over all genes, we derive an expression for the total protein number $N_{p}^{\rm tot}$ in terms of the message cost $\lambda$ and the normalization constant $A$:
@@
This message cost estimate $\hat{\lambda}$ can then be plugged into the metabolic load definition:
$N_{0}\equiv L_{0}+\sum_{i}(\lambda+\varepsilon_{i})\mu_{m,i},$ (108)
to estimate its size:
@@
A.18.1 Detailed protocol
We first estimate the message numbers, as described in Sec. C.2.3, from data. For each gene $i$, we set the optimal message number equal to the observed message number and then compute the optimal overabundance from the message number using Eq. 62. (Since the result is independent of the assumed $\Lambda$ value, we set an arbitrary initial value of $\Lambda=10^{-5}$.) We then use these single gene optimal message number and overabundances to compute $A$ using Eq. 105. In Eqs. 107 and 109, we use the $N^{\rm tot}_{p}$ from Tab. S2. $N_{m}^{\rm tot}$ is computed by summing the estimated message numbers.
A.18.2 Estimate the message cost and metabolic load in yeast
@@
A.18.3 Estimate the message cost and metabolic load in human cells
@@
A.18.4 Estimate of the modified relative load in bacterial cells
@@
$N_{0}$ $=$ $\sum_{i}(\lambda+\varepsilon){\mu_{m,i}}=(\lambda+\varepsilon)N_{m},$ (118)
@@ B.2 Results: Overabundance is a robust prediction
Model organism Growth condition Doubling time: Message lifetime: Message recycling ratio: messages /cell: messages /cell-cycle: proteins: translation efficiency: translation rate:
$T$ $\tau_{m}=\gamma_{m}^{-1}$ $m=T/\tau_{m}$ $N^{\rm tot}_{m/c}$ $N^{\rm tot}_{m}$ $N^{\rm tot}_{p}$ $\varepsilon$ $\beta_{p}$ (h-1)
Escherichia coli LB 30 min [35] $2.5$ min [36] 12 $7.8\times 10^{3}$ [44] $9.4\times 10^{4}$ $3\times 10^{6}$ [45] $22$ $530$
(E. coli) M9 90 min [35] $2.5$ min [36] 36 $2.4\times 10^{3}$ [44] $8.6\times 10^{4}$ $3\times 10^{6}$ [45] $24$ $580$
Sacchromyces cerevisiae (Yeast–haploid) YEPD 90 min [46] 22 min [47] 4 $2.9\times 10^{4}$ [48] $1\times 10^{5}$ $5\times 10^{7}$ [49] $4\times 10^{2}$ $410$
Mus musculus (Mammalian-mouse) Tissue 27.5 h [23] 15 h [23] 1.8 $1.7\times 10^{5}$ [23] $3\times 10^{5}$ [23] $3\times 10^{9}$ [23] $1\times 10^{4}$ $660$
Homo sapiens (Human) Tissue 24 h [50] 14 h [51] 1.7 $3.6\times 10^{5}$ [46] $5\times 10^{5}$ $2\times 10^{9}$ [45] $4\times 10^{3}$ 120
@@ C.1 Methods: Selection of central dogma parameter estimates
Essential gene classification. The classification of essential genes in E. coli comes from the construction of the Keio knockout collection from Baba et al. [53].
@@ C.1.3 Human data
Figure S9: The one message rule. Panel A: One-message-rule for essential genes. For highly transcription genes (high $\mu_{m}$), little compensation for noise is required and the optimal message number tracks with the threshold message number $n_{m}$. However, as the threshold message number approaches one ($n_{m}\rightarrow 1$), the noise is comparable to the mean, and the optimal message number $\mu_{m}$ increases to compensate for the noise. As a result, a lower threshold of roughly one message per cell-cycle is required for essential genes. This threshold is predicted for both fixed (dashed) and optimized translation efficiency (solid). The threshold is weakly dependent on relative load $\Lambda$. Panel B: A one message threshold is observed in three evolutionarily-divergent organisms. As predicted by the RLTO model, essential, but not nonessential genes, are observed to be expressed above a one message per cell-cycle threshold. All organisms have roughly similar distributions of message number for essential genes, which are not observed for message numbers below a couple per cell cycle. Panel C: The distribution of gene transcription rate. The typical transcription rate varies by two orders-of-magnitude between organisms. Panel D: The distribution of gene cellular message number. There is also a two-order-of-magnitude variation between typical cellular message numbers. No consistent lower threshold is observed for either statistic.
@@ C.2.2 Estimating the transcription rate: $\beta_{m}$
To estimate the transcription rate for gene $i$, we start from the estimated cellular message number $N_{m/c,i}$ and use the Paulsson model prediction for the cellular message number:
@@ C.2.3 Estimating the message number: $\mu_{m}$
To estimate the message number of gene $i$, we use the predicted value from the steady-state stochastic kinetic model (see A.1.1):
@@ C.3 Results: Histograms of central dogma transcriptional statistics
Gene Message Annotated function Essential (E)/
name number: from Ecocyc Nonessential (N)
@@
alsK 0.3 The alsK gene encodes a D-allose kinase. Its role in the degradation of D-allose is unclear; AlsK is not required for utilization of a D-allose carbon source; this effect may be due to the presence of other ambiguous sugar kinases within E. coli K-12. E, N, N
bcsB 0.4 BcsB is encoded in a predicted operon together with bcsA, bcsZ and bcsC. In other organisms, these genes are involved in cellulose biosynthesis, a characteristic of the rdar (red, dry and rough) morphotype. However, the K-12 laboratory strain of E. coli does not show a rdar morphotype and does not produce cellulose. E, N, N
entD 0.4 AcpS is the founding member of a 4’-phosphopantetheinyl (P-pant) transferase protein family that includes E. coli EntD, E. coli o195 protein, and Bacillus subtilis Sfp; family members share two conserved motifs but relatively low sequence identity overall. E, N, N
yafF 0.4 No information about this protein was found by a literature search conducted on April 19, 2017. E,-, N
yagG 0.6 yagGH is predicted to be a member of the XylR regulon; its products may mediate transport (YagG) and hydrolysis (YagH) of xylooligosaccharides; putative XylR and CRP binding sites are identified upstream of yagGH. E,-, N
yceQ 0.2 No information about this protein was found by a literature search conducted on July 12, 2017. E, E, N
ydiL 0.2 No information about this protein was found by a literature search conducted on April 7, 2017. E, N, N
yhhQ 0.4 YhhQ is an inner membrane protein implicated in the uptake of queuosine (Q) precursors - 7-cyano-7-deazaguanine (preQ0) and 7-aminomethyl-7-deazaguanine (preQ1) - for Q salvage. Q-modified tRNA is absent in $\Delta$queD and $\Delta$queD $\Delta$yhhQ strains grown in minimal media with glycerol; Q-modified tRNA is detected when a $\Delta$queD strain is grown in minimal media plus 10 nM preQ0 or preQ1 but is absent when a $\Delta$queD $\Delta$yhhQ strain is grown under these conditions. yhhQ expressed from a plasmid restores the presence of Q-modified tRNA in a $\Delta$queD $\Delta$yhhQ strain. E,-, N
yibJ 0.3 No information about this protein was found by a literature search conducted on July 9, 2018. E, N, N
ymfK 0.4 YmfK is a component of the relic lambdoid prophage e14 and is likely the SOS-sensitive repressor. It is similar to the P34 gene of the Shigella flexneri bacteriophage SfV and belongs to the LexA group of SOS-response transcriptional repressors. E, E, E
@@ Appendix D Analysis of gene-expression noise
This section provides a detailed development of gene expression noise. We continue the discussion of the model from Sec. A.1 that provided a self-contained development of the noise models developed by others which are the input to the RLTO model. Secs. D.1-D.6 describe the RLTO prediction of non-canonical noise scaling and the test of this model.
D.1 Results: RLTO model predicts non-canonical noise scaling
@@
In E. coli, the translation efficiency is roughly constant (i.e.  $\hat{\mu}_{p}\propto\hat{\mu}_{m}$ , Fig. 4D) and therefore Eq. 132 is consistent with the canonical noise model (Eq. 130). However, in eukaryotes, the translation efficiency grows with message number (i.e.  $\hat{\mu}_{p}\propto\hat{\mu}_{m}^{2}$ , Fig. 4BC). If we substitute this proportionality into Eq. 132, we predict the non-canonical noise scaling:
@@
D.2 Methods: Analysis of gene expression noise